Lindstrom Company produces two fountain pen models. Information about its products follows: Sales revenue Less: Variable costComplete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3 Required 4 Determine LindRequired 1 Required 2 Required 3 Required 4 Calculate Lindstroms break-even point in units and in sales revenue. (Round yourRequired 1 Required 2 Required 3 Required 4 Calculate the number of units that Lindstrom must sell to earn a $130,000 profit.Calculate Lindstroms margin of safety and margin of safety as a percentage of sales if it sells 7,500 total pens. (Round you
Answer #1
1 Weighted-average contribution margin:
Contribution margin Total units sold Contribution margin per unit Weight (Based on units sold) Weighted-average contribution margin
a b c=a/b d e=c*d
Product A 92400 5000 18.48 0.5 9.24
Product B 154800 5000 30.96 0.5 15.48
10000 24.72
Weighted-average contribution margin=$ 24.72
Weighted-average contribution margin ratio:
Contribution margin Sales revenue CM Ratio Weight (Based on units sold) Weighted-average contribution margin ratio
a b c=a/b d e=c*d
Product A 92400 165000 0.56 0.5 0.28
Product B 154800 215000 0.72 0.5 0.36
0.64
Weighted-average contribution margin ratio=0.64=64%
2 Break-even point in units=Fixed cost/Weighted-average CM=87500/24.72=3539.644=3540
Break-even point in sales revenue=Fixed cost/Weighted-average CM ratio=87500/64%=$ 136718.8
3 Units to be sold=(Fixed cost+Desired profit)/Weighted-average CM=(87500+130000)/24.72=8798.544=8799  
4 Margin of safety in units=Actual sales-Break-even sales=7500-3540=3960
Margin of safety as percentage of sales=(Actual Sales-Break-even sales)/Actual sales=(7500-3540)/7500=3960/7500=0.528=52.8%